. Two impedances Z1 and Z2 are connected in parallel. The first branch takes a leading current of 16 A and has a resistance of 5 Ω, while the second branch takes a lagging current at 0.8pf. The applied voltage is 100+j200 V and the total power is 5 kW. Find branch impedances, total circuit impedance, branch currents and total circuit current.


1
Expert's answer
2021-03-02T18:10:41-0500

Current in the first branch: "I_1=16\ A"

Applied Voltage: "U=100+j200\ V"

Total Power: "P=5\ kW"


Then:

"P=UI"

"5000 =I (100 + j200)"

"I=5000/(100+j200)=10-j20\ A"

Impedance in the first branch:

"Z_1=U/I_1=(100 + j200)/16=6.25+j12.5 \Omega"


"I_2=I-I_1=10-j20-16=-6-j20\ A"

"Z_2=U/I_2=−10.54+j1.83\ \Omega"

Total impedance:

"Z=U/I=\frac{100+j200}{10-j20}=\frac{(100+j200)(10-j20)}{500}=\frac{5000}{500}=10\ \Omega"